If (ð‘¥, ð‘¦, ð‘§) =ð‘‹ï»¿^ 2 𑌠− 2 ð‘Œï»¿^ 2ð‘ ^3, find ∇𑓠at the point (1, −1,2)


1
Expert's answer
2022-03-21T13:35:15-0400

"\nabla f=\left( \frac{\partial f}{\partial x},\frac{\partial f}{\partial y},\frac{\partial f}{\partial z} \right) =\left( 2xy,x^2-4yz^3,-6y^2z^2 \right) \\
\nabla f(1,-1,2)=(-2,33,-24)"